# Birkhoff–Kellogg type results in product spaces and their application to differential systems

### v1

@@ 3. Some applications to differential systems
We apply now the above results to some classes of systems of BVPs in the context of ODEs and PDEs.
3.1. Eigenvalues for a systems of elliptic PDEs
@@
where $\Omega\subset\mathbb{R}^{n}$ denotes the open unit ball in $\mathbb{R}^{n}$, $f:\overline{\Omega}\times\mathbb{R}_{+}\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ and $g:\overline{\Omega}\times\mathbb{R}_{+}\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ are continuous functions. It is folklore that the system (3.1) can be rewritten in the integral form
$\left\{\begin{array}[]{l}u(x)=\displaystyle\lambda_{1}\int_{\Omega}k(x,y)\,f(y,u(y),v(y))\,dy,\\[8.5359pt] v(x)=\displaystyle\lambda_{2}\int_{\Omega}k(x,y)\,g(y,u(y),v(y))\,dy,\end{array}\right$. (3.2)
where $k$ is the Green’s function associated to the linear PDE
$\Delta u=h(x)\ \text{ in }\Omega,\quad u=0\ \text{ on }\partial\,\Omega,$
with $h$ a continuous function. We make use of the cone of positive functions in the space $\mathcal{C}(\overline{\Omega})$ of continuous functions endowed with the usual supremum norm $\left\|u\right\|_{\infty}=\max_{x\in\overline{\Omega}}\left|u(x)\right|$. Namely, we let
$P:=\{u\in\mathcal{C}(\overline{\Omega}):u\geq 0\}.$
With these ingredients we can state the following Theorem.
@@ 2. $b)$
Let us consider the Banach spaces $X=Y=\mathcal{C}(\overline{\Omega})$ and the cones $K_{1}=K_{2}=P$. By classical PDE theory, the operator
@@
defined by
$\begin{array}[]{l}T_{1}(u,v)(x)=\displaystyle\int_{\Omega}k(x,y)\,f(y,u(y),v(y))\,dy,\\[8.5359pt] T_{2}(u,v)(x)=\displaystyle\int_{\Omega}k(x,y)\,g(y,u(y),v(y))\,dy,\end{array}$ (3.3)
is compact. Now let us consider $(u,v)\in P\times P$ such that $\left\|u\right\|_{\infty}=r_{1},\left\|v\right\|_{\infty}\leq r_{2}$. For every $x\in\overline{\Omega}$ we have
@@
$\inf_{\left\|u\right\|_{\infty}=r_{1},\ \left\|v\right\|_{\infty}\leq r_{2}}\left\|T_{1}(u,v)\right\|_{\infty}\geq\sup_{x\in\overline{\Omega}}\int_{\Omega}k(x,y)\,\underline{f}(y)\,dy>0.$
@@
Now fix $r_{1},r_{2}>0$, then the conditions $a)$ and $b)$ of Theorem 3.1 are satisfied with the choice of $\underline{f}(x,y)=\underline{g}(x,y)\equiv 1$, since a direct calculation gives
$\sup_{(x,y)\in\overline{\Omega}}\int_{\Omega}k((x,y),(w,z))\,d(w,z)=\sup_{(x,y)\in\overline{\Omega}}\dfrac{1}{4}(1-x^{2}-y^{2})=\dfrac{1}{4}.$
@@ 3.2. Eigenvalues for a system of ODEs
$\left\{\begin{array}[]{l}u^{\prime\prime}(t)+\lambda_{1}f(t,u(t),v(t))=0,\ t\in[0,1],\\ v^{\prime\prime}(t)+\lambda_{2}g(t,u(t),v(t))=0,\ t\in[0,1],\\ u(0)=u(1)=0=v^{\prime}(0)=v(1)-\frac{1}{2}v^{\prime}(1),\end{array}\right$. (3.6)
@@
$k_{1}(t,s)=\left\{\begin{array}[]{ll}{(1-t)s},&\quad s\leq t,\\ {t(1-s)},&\quad s>t,\end{array}\right$.
@@
In this case we utilize the space $C[0,1]$, endowed with the usual supremum norm $\|u\|_{\infty}:=\max_{t\in[0,1]}|u(t)|$, and work within the product of the cone of positive functions
$K=\bigl{\{}u\in C[0,1]:u\geq 0,\displaystyle{\min_{t\in[1/4,3/4]}}u(t)\geq\frac{1}{4}\|u\|_{\infty}\bigr{\}}$ (3.8)
with the space itself.
With these ingredients we can state the following Theorem.
@@
Let $r_{1},r_{2}$ be positive constants and suppose that exist two continuous functions $\underline{f},\underline{g}:[0,1]\rightarrow\mathbb{R}_{+}$ such that the following conditions hold:
@@ 1. $a)$
    $f(t,u,v)\geq\underline{f}(t)$ on $1/4,3/4]\times[r_{1}/4,r_{1}]\times[-r_{2},r_{2}]$ and
    $\sup_{t\in[1/4,3/4]}\displaystyle\int_{1/4}^{3/4}k_{1}(t,s)\underline{f}(s)\,ds>0;$
@@ 2. $b)$
    $g(t,u,v)\geq\underline{g}(t)$ on $0,1]\times[0,r_{1}]\times[-r_{2},r_{2}]$ and
@@
Let us consider the Banach spaces $X=Y=C[0,1]$ and the cone $K$ as in (3.8). Note that the operator
@@
$\begin{array}[]{r}T_{1}(u,v)(t):=\displaystyle\int_{0}^{1}k_{1}(t,s)f(s,u(s),v(s))\,ds,\\[8.5359pt] T_{2}(u,v)(t):=\displaystyle\int_{0}^{1}k_{2}(t,s)g(s,u(s),v(s))\,ds.\end{array}$ (3.9)
is compact.
Firstly, let us take $(u,v)\in K\times C[0,1]$ such that $\left\|u\right\|_{\infty}=r_{1},\left\|v\right\|_{\infty}\leq r_{2}$. Note that for every $t\in[1/4,3/4]$ we have $u(t)\geq r_{1}/4$ and, furthermore, we have
$\|T_{1}(u,v)\|_{\infty}\geq T_{1}(u,v)(t)=\displaystyle\int_{0}^{1}k_{1}(t,s)f(s,u(s),v(s))\,ds\\ \geq\displaystyle\int_{1/4}^{3/4}k_{1}(t,s)f(s,u(s),v(s))\,ds\geq\displaystyle\int_{1/4}^{3/4}k_{1}(t,s)\underline{f}(s)\,ds.$
Then we get
$\|T_{1}(u,v)\|_{\infty}\geq\sup_{t\in[1/4,3/4]}\displaystyle\int_{1/4}^{3/4}k_{1}(t,s)\underline{f}(s)\,ds.$ (3.10)
@@
$\inf_{\left\|u\right\|_{\infty}=r_{1},\ \left\|v\right\|_{\infty}\leq r_{2}}\left\|T_{1}(u,v)\right\|_{\infty}\geq\sup_{t\in[1/4,3/4]}\displaystyle\int_{1/4}^{3/4}k_{1}(t,s)\underline{f}(s)\,ds>0.$
Secondly, let us take $(u,v)\in K\times C[0,1]$ such that $\left\|u\right\|_{\infty}\leq r_{1},\left\|v\right\|_{\infty}=r_{2}$. Note that
@@
$\inf_{\left\|u\right\|_{\infty}\leq r_{1},\ \left\|v\right\|_{\infty}=r_{2}}\left\|T_{2}(u,v)\right\|_{\infty}\geq\displaystyle\int_{0}^{1}-k_{2}(1,s)\underline{g}(s)\,ds=\frac{1}{2}\displaystyle\int_{0}^{1}\underline{g}(s)\,ds>0.$
@@
$\left\{\begin{array}[]{l}u^{\prime\prime}(t)+\lambda_{1}t(1+u^{2}v^{2})=0,\ t\in[0,1],\\ v^{\prime\prime}(t)+\lambda_{2}te^{uv}=0,\ t\in[0,1],\\ u(0)=u(1)=0=v^{\prime}(0)=v(1)-\frac{1}{2}v^{\prime}(1).\end{array}\right$. (3.12)
Now fix $r_{1},r_{2}>0$, then the choice of $\underline{f}(t)\equiv\frac{1}{4}$ gives, by direct calculation,
$\sup_{t\in[1/4,3/4]}\displaystyle\int_{1/4}^{3/4}k_{1}(t,s)\frac{1}{4}\,ds=\frac{3}{128}>0,$
@@
Then the conditions $a)$ and $b)$ of Theorem 3.3 are satisfied. Furthermore since the pair $(r_{1},r_{2})$ can be chosen arbitrarily in $(0,+\infty)^{2},$ we obtain the existence of two distinct families of uncountably many pairs: both $(\lambda_{1,1},\lambda_{2,1})$, with $\lambda_{i,1}>0$, $i=1,2$, and $(\lambda_{1,2},\lambda_{2,2})$, with $\lambda_{1,2}>0$ and $\lambda_{2,2}<0$, each of them with the associated eigenfunctions of prescribed norms.
@@ Acknowledgements
A. Calamai and G. Infante are members of the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM). G. Infante is a member of the UMI Group TAA and the “The Research ITalian network on Approximation (RITA)”. J. Rodríguez–López has been partially supported by the VIS Program of the University of Calabria, by Ministerio de Ciencia y Tecnología (Spain), AEI and Feder, grant PID2020-113275GB-I00, and by Xunta de Galicia, grant ED431C 2023/12. This study was partly funded by: Research project of MIUR (Italian Ministry of Education, University and Research) Prin 2022 “Nonlinear differential problems with applications to real phenomena” (Grant Number: 2022ZXZTN2).

### v2

@@ 3. Some applications to differential systems
In this Section we present some applications of the above results to the context of systems of ODEs and PDEs.
3.1. Eigenvalues for a system of elliptic PDEs
@@
where $\Omega\subset\mathbb{R}^{n}$ denotes the open unit ball in $\mathbb{R}^{n}$, $f:\overline{\Omega}\times\mathbb{R}_{+}\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ and $g:\overline{\Omega}\times\mathbb{R}_{+}\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$ are continuous functions.
It is folklore (see for example [19, Section 7.2] or [11, Section 4.2]), that if $h$ is sufficiently smooth, the unique solution of the linear PDE
$\Delta w=h(x)\ \text{ in }\Omega,\quad w=0\ \text{ on }\partial\,\Omega,$ (3.2)
can be represented in the integral form
$w(x)=\displaystyle\int_{\Omega}k(x,y)\,h(y)\,dy=:L(h)(x),$ (3.3)
where $k$ is the Green’s function associated to the PDE. The explicit formula for the Geeen’s function for the ball depends on the dimension of the space and can be found for example in [11, Section 2]. It is also well known that the Green’s function is non-negative and continuous for $x\neq y$ and singular on the diagonal. The recent paper [29] provides the expression of the Green’s function and studies its properties in more general domains. Note that if the function $h$ that occurs in (3.2) is continuous one may still study the solutions of the corresponding Hammerstein integral equation (3.3) and may interpret those as weak solutions of (3.2), see [19, Section 7.2]. Now, denote by
$P:=\{u\in\mathcal{C}(\overline{\Omega}):u\geq 0\},$
where $\mathcal{C}(\overline{\Omega})$ is the space of continuous functions endowed with the usual supremum norm $\left\|u\right\|_{\infty}=\max_{x\in\overline{\Omega}}\left|u(x)\right|$. Note that $P$ is a cone in $\mathcal{C}(\overline{\Omega})$ and the linear operator $L$ leaves the cone $P$ invariant and is compact, see for example [19, Section 7.2] and [29, Proposition 5.1].
With these ingredients in mind, we say that a pair $(u,v)\in\mathcal{C}(\overline{\Omega})\times\mathcal{C}(\overline{\Omega})$ is a (weak) solution of problem (3.1) if $(u,v)$ is a solution of the integral system
$\left\{\begin{array}[]{l}u(x)=\displaystyle\lambda_{1}\int_{\Omega}k(x,y)\,f(y,u(y),v(y))\,dy,\\[8.5359pt] v(x)=\displaystyle\lambda_{2}\int_{\Omega}k(x,y)\,g(y,u(y),v(y))\,dy.\end{array}\right$.
In other words, $(u,v)$ is a solution of (3.1) if and only if
$\left\{\begin{array}[]{l}u=\displaystyle\lambda_{1}T_{1}(u,v),\\[8.5359pt] v=\displaystyle\lambda_{2}T_{2}(u,v),\end{array}\right$. (3.4)
where
$\begin{array}[]{l}T_{1}(u,v)(x)=\displaystyle\int_{\Omega}k(x,y)\,f(y,u(y),v(y))\,dy,\\[8.5359pt] T_{2}(u,v)(x)=\displaystyle\int_{\Omega}k(x,y)\,g(y,u(y),v(y))\,dy.\end{array}$
We are now in position to state the following Theorem.
@@ 2. $b)$
Let us consider the Banach spaces $X_{1}=X_{2}=\mathcal{C}(\overline{\Omega})$ and the cones $K_{1}=K_{2}=P$. Due to the compactness and invariance properties of the operator $L$ defined in (3.3), combined with the continuity of $f,g$, the operator
@@
defined in (3.1), is compact. Now let us consider $(u,v)\in P\times P$ such that $\left\|u\right\|_{\infty}=r_{1},\left\|v\right\|_{\infty}=r_{2}$. For every $x\in\overline{\Omega}$ we have
@@
$\inf_{\left\|u\right\|_{\infty}=r_{1},\ \left\|v\right\|_{\infty}=r_{2}}\left\|T_{1}(u,v)\right\|_{\infty}\geq\sup_{x\in\overline{\Omega}}\int_{\Omega}k(x,y)\,\underline{f}(y)\,dy>0.$
@@
Now, note that the function
$w(x,y):=\dfrac{1}{4}(1-x^{2}-y^{2}),$
is the unique solution of the “torsion problem”
$\Delta w=1\ \text{in}\ \Omega,\quad w=0\ \text{on}\ \partial\,\Omega.$
Now fix $r_{1},r_{2}>0$ and observe that the conditions $a)$ and $b)$ of Theorem 3.1 are satisfied with the choice of $\underline{f}(x,y)=\underline{g}(x,y)\equiv 1$, since
$\sup_{(x,y)\in\overline{\Omega}}\int_{\Omega}k((x,y),(\xi,\eta))\,d(\xi,\eta)=\sup_{(x,y)\in\overline{\Omega}}\dfrac{1}{4}(1-x^{2}-y^{2})=\dfrac{1}{4}.$
@@ 3.2. Eigenvalues for a system of ODEs
$\left\{\begin{array}[]{ll}u^{\prime\prime}(t)+\lambda_{1}f(t,u(t),v(t))=0,&t\in[0,1],\\ v^{\prime\prime}(t)+\lambda_{2}g(t,u(t),v(t))=0,&t\in[0,1],\\ u^{\prime}(0)=u(1)+u^{\prime}(1)=0,\\ v^{\prime}(0)=v(1)-\frac{1}{2}v^{\prime}(1)=0,\end{array}\right$. (3.7)
@@
$k_{1}(t,s)=\left\{\begin{array}[]{ll}{2-t},&\quad s\leq t,\\ {2-s},&\quad s>t,\end{array}\right$.
@@
We point that the kernel $k_{1}$ is continuous, non-negative, and has been studied in details in [21, 23] while the kernel $k_{2}$ is continuous, changes sign, and can be obtained via a direct calculation. In this case we utilize the space $C[0,1]$, endowed with the usual supremum norm $\|u\|_{\infty}:=\max_{t\in[0,1]}|u(t)|$. We construct, following the computations in [23], the cone
$K=\left\{u\in C[0,1]:\displaystyle{\min_{t\in[0,1]}}u(t)\geq\Bigl(1-\frac{t}{2}\Bigr)\|u\|_{\infty}\right\},$ (3.9)
and consider its product with the space itself.
With these ingredients we can state the following Theorem; the construction of the minorant $\underline{f}$ follows the argument in [22].
@@
Let $r_{1},r_{2}$ be positive constants and suppose that exist two measurable functions $\underline{f}:[0,1]\rightarrow\mathbb{R}_{+}$ and $\underline{g}:[0,1]\rightarrow\mathbb{R}_{+}$ such that the following conditions hold:
@@ 1. $a)$
    $f(t,u,v)\geq\underline{f}(t)$ on $0,1]\times[r_{1}/2,r_{1}]\times[-r_{2},r_{2}]$ and
    $\sup_{t\in[0,1]}\displaystyle\int_{0}^{1}k_{1}(t,s)\,\underline{f}(s)\,ds>0;$
@@ 2. $b)$
    $g(t,u,v)\geq\underline{g}(t)$ on $0,1]\times[r_{1}/2,r_{1}]\times[-r_{2},r_{2}]$ and
@@
Let us consider the Banach spaces $X_{1}=X_{2}=C[0,1]$ and the cone $K$ as in (3.9). Note that the operator
@@
$\begin{array}[]{r}T_{1}(u,v)(t):=\displaystyle\int_{0}^{1}k_{1}(t,s)\,f(s,u(s),v(s))\,ds,\\[8.5359pt] T_{2}(u,v)(t):=\displaystyle\int_{0}^{1}k_{2}(t,s)\,g(s,u(s),v(s))\,ds,\end{array}$
is compact, as a consequence of the classical Arzelà–Ascoli Theorem.
Firstly, let us take $(u,v)\in K\times C[0,1]$ such that $\left\|u\right\|_{\infty}=r_{1},\left\|v\right\|_{\infty}=r_{2}$. Note that for every $t\in[0,1]$ we have
$u(t)\geq\Bigl(1-\frac{t}{2}\Bigr)r_{1}\geq\frac{r_{1}}{2}.$
Therefore, we have
$\|T_{1}(u,v)\|_{\infty}\geq T_{1}(u,v)(t)=\displaystyle\int_{0}^{1}k_{1}(t,s)\,f(s,u(s),v(s))\,ds\geq\int_{0}^{1}k_{1}(t,s)\,\underline{f}(s)\,ds,$
for every $t\in[0,1]$. Thus we obtain
$\|T_{1}(u,v)\|_{\infty}\geq\sup_{t\in[0,1]}\displaystyle\int_{0}^{1}k_{1}(t,s)\,\underline{f}(s)\,ds.$ (3.10)
@@
$\inf_{\left\|u\right\|_{\infty}=r_{1},\ \left\|v\right\|_{\infty}=r_{2}}\left\|T_{1}(u,v)\right\|_{\infty}\geq\sup_{t\in[0,1]}\displaystyle\int_{0}^{1}k_{1}(t,s)\,\underline{f}(s)\,ds>0.$
Secondly, let us take $(u,v)\in K\times C[0,1]$ such that $\left\|u\right\|_{\infty}=r_{1},\left\|v\right\|_{\infty}=r_{2}$. Note that
@@
$\inf_{\left\|u\right\|_{\infty}=r_{1},\ \left\|v\right\|_{\infty}=r_{2}}\left\|T_{2}(u,v)\right\|_{\infty}\geq\displaystyle\int_{0}^{1}-k_{2}(1,s)\,\underline{g}(s)\,ds=\frac{1}{2}\displaystyle\int_{0}^{1}\underline{g}(s)\,ds>0.$
@@
$\left\{\begin{array}[]{l}u^{\prime\prime}(t)+\lambda_{1}t(u^{2}+v^{2})=0,\ t\in[0,1],\\ v^{\prime\prime}(t)+\lambda_{2}te^{uv}=0,\ t\in[0,1],\\ u^{\prime}(0)=u(1)+u^{\prime}(1)=0=v^{\prime}(0)=v(1)-\frac{1}{2}v^{\prime}(1).\end{array}\right$.
Now fix $r_{1},r_{2}>0$, then the choice of $\underline{f}(t)=\dfrac{t\,r^{2}_{1}}{4}$ gives, by direct calculation,
$\sup_{t\in[0,1]}\displaystyle\int_{0}^{1}k_{1}(t,s)\frac{t\,r^{2}_{1}}{4}\,ds=\frac{r^{2}_{1}}{4}>0,$
@@
Then the conditions $a)$ and $b)$ of Theorem 3.3 are satisfied. Furthermore since the pair $(r_{1},r_{2})$ can be chosen arbitrarily in $(0,+\infty)\times(0,+\infty)$, we obtain the existence of two distinct families of uncountably many pairs: both $(\lambda_{1,1},\lambda_{2,1})$, with $\lambda_{i,1}>0$, $i=1,2$, and $(\lambda_{1,2},\lambda_{2,2})$, with $\lambda_{1,2}>0$ and $\lambda_{2,2}<0$, each of them with the associated eigenfunctions of prescribed norms.
@@ Acknowledgements
The authors would like to thank the Referee the careful reading of the manuscript and the constructive comments. A. Calamai and G. Infante are members of the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM). G. Infante is a member of the UMI Group TAA and the “The Research ITalian network on Approximation (RITA)”.
Funding
J. Rodríguez–López has been partially supported by the VIS Program of the University of Calabria, by Ministerio de Ciencia y Tecnología (Spain), AEI and Feder, grant PID2020-113275GB-I00, and by Xunta de Galicia, grant ED431C 2023/12. This study was partly funded by: Research project of MIUR (Italian Ministry of Education, University and Research) Prin 2022 “Nonlinear differential problems with applications to real phenomena” (Grant Number: 2022ZXZTN2).
Conflicts of interest
The authors declare no conflict of interest.
Contribution statement
All authors contributed equally to this manuscript.